OpenStudy (anonymous):

State the distance formula . State the midpoint formula. Help me please I don't understand this question well both.

OpenStudy (lilcathyx):

distance formula =square root (x2-x1)^2 + (y2-y1)^2

hero (hero):

\[D =\sqrt{\left(x_2 - x_1\right)^2 + \left(y_2 - y_1\right)^2}\]

hero (hero):

\[M = \left(\frac{x_1 + x_2}{2}, \frac{y_1 + y_2}{2}\right)\]